Pick 3 cards. What is the probability that they are all the same suit?
This is how I attempted to answer this question:
P(same suit) = (13/52)x(12/51)x(11/50) = 1716/132600 which reduced down to 143/11050
I am unsure if I went about this the right way and wonder what the right way to go about it is?
I would appreciate any help you could offer regarding this.

Answer by richwmiller(17219)

You can reduce a lot more.
13/52=1/4
1/4*12/51*11/50
12/50=6/25
1/4*6/25*11/51
6/4=3/2
3*11/(2*25*51)
3/2*51=1/34
1/34*11/25
11/(34*25)
11/850
.0129
1.29%